# 1st Dimitrios Bouzoulas Energy and Mechanical Engineering

### v1

@@ Abstract
Pedestrian detection in RGB images is a key task in pedestrian safety, as the most common sensor in autonomous vehicles and advanced driver assistance systems is the RGB camera. A challenge in RGB pedestrian detection, that does not appear to have large public datasets, is low-light conditions. As a solution, in this research, we propose an automated infrared-RGB labeling pipeline. The proposed pipeline consists of 1) Infrared detection, where a fine-tuned model for infrared pedestrian detection is used 2) Label transfer process from the infrared detections to their RGB counterparts 3) Training object detection models using the generated labels for low-light RGB pedestrian detection. The research was performed using the KAIST dataset. For the evaluation, object detection models were trained on the generated autolabels and ground truth labels. When compared on a previously unseen image sequence, the results showed that the models trained on generated labels outperformed the ones trained on ground-truth labels in 6 out of 9 cases for the mAP@50 and mAP@50-95 metrics. The source code for this research is available at https://github.com/BouzoulasDimitrios/IR-RGB-Automated-LowLight-Pedestrian-Labeling
Index Terms:
Pedestrian Detection, Low-Light Conditions.
@@ I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian related traffic accidents [1]. According to the World Health Organization (WHO) [2], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [3] on the data provided by the transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
@@
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in edge case conditions such as low-light conditions as RGB model accuracy is highly affected by lighting [4]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared(IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal we leverage the lighting invariance of infrared cameras as the visibility of infrared light is not affected by the lack of visible light. In our approach, we utilise the benefits of infrared images to detect pedestrians in infrared and then transfer the labels to the corresponding RGB image pair. Our proposed approach relies on aligned infrared and RGB images, which can be conveniently collected with a research vehicle equipped with suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work.
The proposed pipeline shows a range of results. As a method the proposed pipeline has not been previously proposed for automated labelling. We allow the reuse and further development by open-sourcing the implementation. The experimental results show that the labels generated from the pipeline can match and in many cases outperform ground truth label performance.
@@ II-A Pedestrian Detection
Aiming to enable AVs and ADAS to detect pedestrians reliably, multiple approaches using various sensors and sensor combinations have been proposed. Multiple camera-based solutions have been proposed over the years [5], targeting use cases with favourable lighting conditions. A recent advancement in purely camera-based pedestrian detection is [6], where a modified YOLOv5 model was proposed resulting in mAP@50 76.9% and mAP@50-95 48.3% mAP scores increase on the WiderPerson dataset [7] and a reduced number of floating point operations making the architecture more suitable for real-time applications. In a similar approach [8] proposed a modified YOLOv8 network achieving an mAP@50 of 54.7% mAP@50-95 of 32.6% and a lower number of floating point operations on the WiderPerson dataset. In [9], an infrared camera was utilised for pedestrian detection. Saliency maps were extracted from infrared images and used as an attention mechanism. The approach was used for enhanced pedestrian detection. The results on the KAIST dataset showed a lowered miss rate of 13.4% for daytime and 19.3% for nighttime compared to the baseline in day and night.
For low-light pedestrian detection, more diverse approaches have been proposed to solve the challenges posed by the poor lighting conditions. To achieve improved focus on pedestrian features [10] proposed a fore-background attention mechanism that enhances pedestrian features while reducing background noise, resulting in improved low-light pedestrian detection. The proposed method scored accurate results across multiple datasets compared to the baseline and other state-of-the-art nighttime pedestrian detection methods.
Sensor fusion approaches are a common solution for pedestrian detection, as they offer increased robustness to a wide variety of conditions, including low-light scenarios. Common sensor fusion-based approaches utilize a RGB-Lidar sensor combination. Following the RGB-Lidar approach [11] proposed a two-branch multilevel pedestrian detection network, and when tested on the KITTI dataset [12] achieved an mAP0.5 of 91.6% and 26.7 fps on the Jetson Orin, making the application both accurate and efficient for real-time applications. Similarly, in [13], using the RGB-Lidar sensor combination and applying a stable diffusion-based approach an improvement of the mAP0.5 metric was achieved of 4.2% for BEVDepth and 4.6% for BEVStereo on the nuScenes dataset.
Similarly to the RGB-Lidar methods a wide range of approaches has been suggested for RGB-Infrared/Thermal sensor combination. With [14] proposing a dual network architecture based on YOLOv5, tested on a wide range of publicly available datasets and achieving mAP@0.5 of 96.6%-INRIA, 89.2% CVC09, 90.5% LSIFIR, 56% FLIR-ADAS, 79.8% CVC14, 72.3% Nightowls, 53.3%-KAIST. Using an SSD-like halfway fusion model network was proposed in [15], where features were extracted from thermal and RGB images with two separate networks before being fused, yielding results for KAIST LLVIP and CVC-14, exceeding state-of-the-art performance for mAP metrcis.
II-B Automatic Labelling
Targeting domain adaptation across different datasets, [16] proposed the usage of pseudo-labels for training pedestrian detection models. The approach used an initial dual-input model trained on a baseline dataset using both RGB and thermal images. This was followed by generating and filtering pseudo labels based on confidence on the target dataset using the pretrained model. The model is then trained on the newly annotated dataset and the entire process is repeated until the model performance becomes stable. The proposed method yielded an All-day mAP@50 score of 80.23% while the same model trained on ground truth labels yielded a mAP@50 86.90% resulting in a performance loss of 6.67% compared to ground truth labels in performance while being fully automated.
To achieve automatic labelling across domains [17] proposed an unsupervised learning transfer framework. The proposed method consists of training a model to perform pedestrian detection, using illumination awareness to generate new data, consequently adapting the network to the new domain, and finally performing the automated labelling task on the new domain. The training and autolabelling happens with three models, a daytime using RGB, a nighttime using thermal and a multimodal one using RGB and thermal. The kept labels are decided by an Illumination estimation network as it estimates which autolabeler performs best based on illumination. This approach achieved the lowest miss rate compared to other methods and state-of-the-art performance on the KAIST and FLIR datasets.
@@ III-A Architecture
First, RGB-IR image pair is taken as input, and pedestrians are detected from the IR image. Then the detections are transferred from the IR frame to the corresponding RGB frame to act as a label. Finally, an RGB prediction model for low-light conditions is trained with the autolabeled data.
@@ III-A1 IR detection
In low-light conditions, pedestrians are usually easier to detect from IR images compared to RGB images. IR images capture the IR radiation emitted from the pedestrians, and the amount of IR radiation emitted is not affected by lighting conditions. As a result, IR images are invariant to lighting conditions, and an IR model trained on daytime data generalizes well to low-light conditions. Additionally, due to the heating effect of the sun, the environment usually emits more IR radiation during the day than at night. Thus, the pedestrian IR footprint stands out from the background better at night. From these premises, we train an IR pedestrian detection model with daytime IR data and use it to detect pedestrians in unseen IR images from low-light conditions. These IR detections are then used as labels for the corresponding RGB frames
@@ III-A2 Label transfer
To train an RGB detection model, the labels must be transferred from the IR frame to the RGB frame. Accurate label transfer requires synchronization and known calibration between the IR and RGB frames. Given a pixel in the IR frame, we need to be able to define its position in the RGB frame. To achieve this, the camera matrices must be known, and the principal points of the IR and RGB cameras must be aligned, or at least one of the cameras must also include depth.
In the KAIST dataset used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
@@ III-B Data
[…]
@@ III-C Evaluation
We chose the YOLOv11[18] for the IR detection task, and trained the m, l, and x variants to compare performance across different model sizes. These are referred to as IR-YOLOv11m, IR-YOLOv11l, and IR-YOLOv11x, respectively. We started with the COCO pretrained weights and fine-tuned the weights for the IR pedestrian detection task using the manually labeled daytime IR images. Training was conducted with the default YOLOv11 hyperparameters for 10 epochs. Labels were then automatically generated for the low-light train samples with each variation. The manual ground truth labels were used as a baseline. Altogether, four different training labels were used in the evaluation:
[…]
We employed YOLOv11 also for the RGB detection task, and again trained the m, l, and x variants. These are referred to as YOLOv11m, YOLOv11l, and YOLOv11,x respectively. We started with the COCO pretrained weights and trained for 10 epochs with the default YOLOv11 hyperparameters. Each variant was trained with each of the training labels, yielding altogether 12 different training label-model combinations.
@@ IV Results
The proposed autolabels were evaluated by comparing the performance of the models trained on autolabels to the models trained on ground truth labels (Table I). Precision, Recall, mAP@50, and mAP@50-95 were chosen as evaluation metrics. The evaluation function presented in [19] was used here. Additionally, examples of detection results are shown in Figure 3.
Table I: Validation test results comparison with green for results above the baseline and red for results below baseline. The precision and recall values have been reported with a confidence threshold of 0.1 and IoU threshold of 0.5.
Labels P R mAP@50 mAP@50-95
Base Model YOLOv11m
IR-YOLOv11m (our) 0.813 0.282 0.306 0.284
IR-YOLOv11l (our) 0.754 0.299 0.311 0.288
IR-YOLOv11x (our) 0.926 0.204 0.313 0.280
KAIST/Ground Truth 0.824 0.269 0.290 0.271
Base Model YOLOv11l
IR-YOLOv11m (our) 0.817 0.306 0.339 0.318
IR-YOLOv11l (our) 0.788 0.325 0.359 0.337
IR-YOLOv11x (our) 0.919 0.204 0.308 0.280
KAIST/Ground Truth 0.822 0.289 0.328 0.306
Base Model YOLOv11x
IR-YOLOv11m (our) 0.894 0.266 0.317 0.294
IR-YOLOv11l (our) 0.896 0.272 0.341 0.317
IR-YOLOv11x (our) 0.935 0.176 0.291 0.265
KAIST/Ground Truth 0.802 0.281 0.318 0.295
In both the mAP@50 and mAP@50-95 metrics, the models trained with the proposed autolabels outperformed the models trained with the ground truth in 6 out of 9 cases. For all prediction model sizes, the highest performance was achieved with some version of our autolabels. Overall, the highest mAP@50 and mAP@50-95 was achieved by YOLOv11l trained on labels generated with IR-YOLOv11l.
The functionality of the proposed method is further verified by the Precision-Recall curves shown in Figure 2. Precision-Recall Curves are presented for YOLOv11m, l and x for generated and ground truth labels with IoU threshold of 0.5. The Precision-Recall curves further verify that the proposed labels match the ground labels in performance and in many cases exceed it as in all cases the Precision-Recall curves follow the same pattern, dictating similar precision recall tradeoffs.
Figure 2: Precission-Recall Curves for YOLOv11m, l and x for Generated and Ground Truth Labels with IoU threshold of 0.5.
@@
IR-YOLOv11m IR-YOLOv11l IR-YOLOv11x GT labels GT
Prediction model YOLOv11m Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
YOLOv11l Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
YOLOv11x Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: Examples of predictions made with different prediction model - training label combinations. Columns 1-3 correspond to the models trained with our autolabels, column 4 to the models trained with ground truth labels and column 5 is the ground truth.

### v2

# 1st Dimitrios Bouzoulas    2nd Eerik Alamikkotervo    3rd Risto Ojala

@@ Abstract
Pedestrian detection in RGB images is a key task in pedestrian safety, as the most common sensor in autonomous vehicles and advanced driver assistance systems is the RGB camera. A challenge in RGB pedestrian detection, which does not appear to have many public datasets or efficient autolabelling pipelines, is low-light conditions. As a solution, in this research, we propose an automated infrared-RGB labelling pipeline. The proposed pipeline consists of 1) Infrared detection, where a fine-tuned model for infrared pedestrian detection is used 2) Label transfer process from the infrared detections to their RGB counterparts 3) Training object detection models using the generated labels for low-light RGB pedestrian detection. The research was performed using the KAIST dataset. For evaluation, three object detection models, DETR, YOLO, and RCNN, were trained on the generated autolabels and ground truth labels. When compared on a previously unseen image sequence, the results showed that the models trained on generated labels outperformed the ones trained on ground-truth labels in 5 out of 6 cases for the mAP@50 and LAMR metrics, and outperformed ground-truth on mAP@50-95 in all cases. Acquired results indicate that the proposed autolabelling pipeline could be used for scalable annotation of low-light datasets for pedestrian detection. The source code for this research is available at https://github.com/BouzoulasDimitrios/IR-RGB-autoamed-low-light-pedestrian-labelling
@@ I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian-related traffic accidents [1]. According to the World Health Organization (WHO) [2], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors, such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions, nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [3] on the data provided by the transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
@@
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in low-light conditions, as RGB model accuracy is highly affected by lighting [4]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared (IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal, we leverage the lighting invariance of infrared cameras as visible light has minimal effect on infrared radiation.
In our approach, we utilise the benefits of infrared images to detect pedestrians in infrared and then transfer the labels to the corresponding RGB image. Our proposed approach relies on aligned infrared and RGB images, which can be conveniently collected with a research vehicle equipped with a suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work.
The proposed pipeline has not previously been implemented for autolabelling tasks and shows a range of positive initial results. The autolabels generated using the proposed pipeline performed on par or outperformed the ground truth labels in many cases in key metrics such as map@50, map@50-95, and LAMR.
@@ II-A Pedestrian Detection
Early pedestrian detection approaches utilize hand-crafted features such as Histogram of Oriented Gradients (HOG) [5]. However, currently, the best performance is achieved by deep neural networks that can learn complex detection mechanisms from the data. Originally, convolutional neural network architectures were proposed but recently also transformer-based architectures have gained popularity in the pedestrian detection task, DETR [6] being a popular example. Convolutional architectures are typically further divided into one-stage and two-stage detectors. One-stage detectors predict the bounding box coordinates and related class probabilities in a single forward pass. Common example is the YOLO family [7]. Two-stage detectors first propose regions of interest and then classify and refine the proposals. Common examples are RCNN [8] and its predecessors Fast RCNN [9] and Faster RCNN [10]. To support the development of the pedestrian detection models variety of datasets have been presented over the years, e.g. Caltech Pedestrian Dataset [11], CityPersons dataset [12], KITTI [13] and WiderPerson [14].
II-B Detection in low-light conditions
In low-light conditions, pedestrian detection from conventional visible-spectrum RGB images is significantly more challenging as the lack of visible light leads to low contrast, loss of color information, and severe noise effects. To improve low-light RGB detection, image enhancement methods which aim to restore the low-light image before feeding it to a standard detector, have been proposed [15, 16]. Improved performance can be achieved by optimizing the enhancement for the object detection task specifically by end-to-end training [17, 18]. However, superior performance can be achieved by leveraging additional sensors like a thermal camera or lidar that provide measurements which are much less affected by the lighting conditions. Lidar-camera fusion for object detection has been extensively researched and [19, 20] are notable examples including low-light conditions in the experiments. RGB-thermal fusion has also been explored in multiple studies. A popular baseline solution is provided in [21] and recent work presenting a more advanced fusion strategy in [22]. In [23, 24] teacher-student architecture is proposed where the teacher has access to RGB and thermal images while the students can only rely on RGB images to train a reliable RGB detector for low-light conditions. While most of the available data is from daytime, some datasets addressing the low-light conditions have been published. KAIST [21], LLVIP [25], FLIR-ADAS [26] and CVC-14 [27] provide RGB and thermal images, and Nightowls [28] only provides RGB images.
II-C Automated learning
Detection models perform poorly outside their training domain, and manual annotation is one of the main bottlenecks for training data generation. Thus, research efforts have also focused on ways to reduce or remove the need for manual annotations. A common approach is unsupervised domain adaptation, where the model is adapted from a source domain to a target domain and annotations are only available in the source domain. In [29] RGB-thermal pedestrian detection model was adapted from the KAIST dataset domain to the CVC-14 dataset domain in unsupervised manner with performance close to supervised methods. Also in [30] unsupervised domain adaptation yields comparable results to supervised methods when moving between the CVC-14, KAIST and FLIR ADAS datasets. Another way to reduce the need for labeling is to transfer the labels between sensors. If reliable detections are available in one sensor frame they can be moved to other sensor frames. In [31] bounding boxes from pre-trained RGB detector were transfered to the corresponding thermal frames to train thermal detection models without the need for manual labeling. Arguably for low-light pedestrian detection transfering the labels in opposite direction would be lucrative. Reliable detections on thermal images could be transferred to the corresponding RGB frames to train a RGB prediction model.
@@ III-A Architecture
First, an RGB-IR image pair is taken as input, and pedestrians are detected from the IR image. Then the detections are transferred from the IR frame to the corresponding RGB frame to act as a label. Finally, an RGB prediction model for low-light conditions is trained with the autolabeled data.
@@ III-A1 IR detection
In low-light conditions, pedestrians are usually easier to detect from IR images compared to RGB images. IR images capture the IR radiation emitted from the pedestrians, and the amount of IR radiation emitted by pedestrians is roughly invariant to visible light conditions. As a result, an IR model trained on daytime data generalizes well to low-light conditions. Additionally, due to the heating effect of the sun, the environment usually emits more IR radiation during the day than at night. Thus, the pedestrian IR footprint stands out from the background better at night. From these premises, we train an IR pedestrian detection model with daytime IR data and use it to detect pedestrians in unseen IR images from low-light conditions. These IR detections are then used as labels for the corresponding RGB frames
@@ III-A2 Label transfer
To train an RGB detection model, the labels must be transferred from the IR frame to the RGB frame. Accurate label transfer requires synchronisation and known calibration between the IR and RGB frames. Given a pixel in the IR frame, we need to be able to define its position in the RGB frame. To achieve this, the camera matrices must be known, and the principal points of the IR and RGB cameras must be aligned, or at least one of the cameras must also include depth.
In the KAIST dataset [21] used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
@@ III-B Data
The proposed autolabelling method is evaluated with the KAIST dataset. KAIST includes $\sim$95k samples in total, where each sample includes an aligned IR-RGB frame pair and manual ground truth annotation for the cyclist, person, and people classes. The people class is used if individual persons cannot be distinguished from a group of people. In our experiments, we excluded the samples that include cyclist or people labels, yielding $\sim$91.8k samples. This exclusion was done due to the three classes being too closely related to each other, resulting in extremely poor performance for any models trained on the RGB low-light portion of the dataset. Of the remaining images, $\sim$52.7k samples are from daytime (sequences 00, 01, 02, 06, 07, 08) and $\sim$23.7k samples from lowlight conditions (sequences 03, 04, 05, 09, 10, 11). The IR models were trained with all of the daytime sequences except sequence 02, which was retained for testing, resulting in $\sim$51k training images and $\sim$7.3k testing images. The low-light RGB models were trained with all of the low-light sequences except sequence 09, which was retained for testing, resulting in $\sim$26.7k training images and $\sim$3.5k testing images.
@@ III-C Evaluation
We chose to evaluate our pipeline with the YOLOv11[32], RCNN [10], and DETR[33] architectures. The specific versions chosen were YOLOv11l, Faster RCNN, and RF-DETR. All three models were trained on the daytime IR frames for 20 epochs with the default hyperparameters. These fine-tuned versions of the models are referred to as IR-YOLO, IR-RCNN, and IR-DETR. For all three models, the models were initialized with COCO [34] pretrained weights. The fine-tuned IR models were used for autolabelling the low-light portion of the dataset. To evaluate how the label confidence threshold affects the final results, two label sets were generated with each model, using 0.25 and 0.5 confidence thresholds. Additionally, the models were also trained with ground truth labels, allowing for performance comparisons between models trained on generated and ground truth labels.
Table I: Number of generated autolabels by each IR pedestrian detection model with different confidence thresholds.
Model Confidence Label Count
IR-YOLO 0.25 26712
IR-YOLO 0.5 16388
IR-DETR 0.25 49280
IR-DETR 0.5 29291
IR-RCNN 0.25 21116
IR-RCNN 0.5 19124
Ground Truth - 26779
Same YOLO, DETR, and RCNN variants were also used for the low-light RGB pedestrian detection. Each of the three models was trained on its IR version labels generated with 0.25 and 0.5 confidence thresholds, and the ground truth labels for comparison. This resulted in a total of 9 low-light RGB pedestrian detection models. Similar to the IR models, each model was trained for 20 epochs with the default hyperparameters using COCO pretrained weights as the starting point.
@@ IV Results
The number of labels generated with each of the IR detection models at 0.25 and 0.5 confidence thresholds are presented in Table I. The KAIST ground truth label count is included for comparison. IR-DETR with 0.25 confidence generated the highest number of labels, and IR-YOLO with 0.5 confidence generated the least amount of labels. To assess the accuracy of the generated labels, standard object detection metrics Precision(P), Recall(R), F1, mAP@50,mAP@50-95 and Log Average Miss Rate (LAMR) are also presented at both confidence levels (Table II). IR-DETR had the highest overall performance on the mAP and LAMR metrics, and the labels generated with IR-DETR at 0.5 confidence had the highest F1 score. YOLO had the second-highest and RCNN the lowest performance on the mAP and LAMR metrics. For RCNN and YOLO 0.25 confidence threshold yielded a higher F1 score than a 0.5 confidence threshold. The precision-recall curves for the IR models are presented in Figure 2.
Table II: Autolabelling accuracy metrics for the IR pedestrian detection models with different confidences. Best result bolded in each column.
Conf. 0.5 Conf. 0.25
Model P R F1 P R F1 mAP@50 mAP@50-95 LAMR $\downarrow$
IR-YOLO 0.836 0.370 0.513 0.746 0.544 0.630 0.635 0.245 0.564
IR-RCNN 0.768 0.391 0.519 0.743 0.435 0.549 0.554 0.208 0.623
IR-DETR 0.788 0.528 0.632 0.545 0.741 0.628 0.688 0.284 0.518
Table III: Validation test results comparison with green for results above the baseline and red for results below baseline. The precision and recall values have been reported with an IoU threshold of 0.5.
Labels P R F1 mAP@50 mAP@50-95 LAMR $\downarrow$
Base Model YOLO
IR-YOLO (Conf. 0.5) 0.941 0.171 0.289 0.290 0.128 0.733
IR-YOLO (Conf. 0.25) 0.924 0.236 0.376 0.321 0.138 0.707
KAIST/Ground Truth 0.902 0.196 0.322 0.261 0.116 0.758
Base Model RCNN
IR-RCNN (Conf. 0.5) 0.955 0.202 0.333 0.290 0.125 0.724
IR-RCNN (Conf. 0.25) 0.944 0.221 0.358 0.307 0.127 0.711
KAIST/Ground Truth 0.963 0.207 0.341 0.292 0.115 0.719
Base Model DETR
IR-DETR (Conf. 0.5) 0.809 0.337 0.475 0.437 0.178 0.635
IR-DETR (Conf. 0.25) 0.722 0.369 0.489 0.442 0.183 0.634
KAIST/Ground Truth 0.835 0.315 0.457 0.397 0.154 0.663
The performance of the low-light RGB models trained with the autolabels is evaluated in Table III. The RGB model trained with the manual ground truth annotation acts as the baseline. The models trained with the proposed autolabels outperformed the models trained with the ground truth in 5 out of 6 cases in mAP@50 and LAMR metrics, and in 6 out of 6 cases in the mAP@50-95 metric. For all models, ground truth label performance in mAP and LAMR metrics was exceeded when using 0.25 label confidence, and for YOLO and DETR, also when using 0.5 label confidence.
The best performance overall was achieved by DETR trained on generated labels with a confidence value of 0.25. The precision-recall curves for each RGB model variant are presented in Figure 3. Each architecture exhibits quite similar precision-recall tradeoffs regardless of the training labels used, while the differences between architectures are significantly larger.
Figure 2: Precision-Recall curves for IR models used for autolabelling.
(a) YOLO
(b) RCNN
(c) DETR
Figure 3: Precision-Recall (PR) curves for the YOLO, RCNN, and DETR RGB detection models when trained on autolabels generated with the IR models versus ground truth labels.
@@
0.25 Conf 0.50 Conf GT labels GT IR GT RGB
Prediction model
DETR
@@
YOLO
@@
RCNN
Figure 4: Examples of predictions made with different prediction model - training label combinations. Columns 1-2 correspond to the models trained with our autolabels, column 3 to the models trained with ground truth labels and columns 4 and 5 are the ground truth in infrared and RGB.